# thanks: G.Y and M.C. contributed equally to this workthanks: G.Y. and M.C. contributed equally to this work

### v1

@@ I Introduction
By defining a target variable, a collection of potential explanatory variables, and a set of available mathematical operators, SR aims to find equations that accurately predict the target variable as a function of the explanatory ones, while remaining concise and interpretable Makke and Chawla (2024); Landajuela et al. (2022); Wang et al. (2023). Various SR frameworks are currently accessible, such as EUREQA Schmidt and Lipson (2009), Operon Burlacu et al. (2020), DSR Petersen et al. (2021), EQL Sahoo et al. (2018), ESR Bartlett et al. (2024a), and AI Feynman Udrescu and Tegmark (2020). These interpretable machine learning tools have recently sparked renewed interest, catalyzing notable progress across both theoretical and applied sciences La Cava et al. (2021); Bernal et al. (2021); Butter et al. (2024); Shao et al. (2022a); Li et al. (2023); Bartlett et al. (2023); Angelis et al. (2023); Desmond et al. (2023); Sousa et al. (2024); Bartlett et al. (2024b, 2023); Thing and Koksbang (2025); Bartlett et al. (2024c); Constantin et al. (2024); Sui et al. (2024); Makke and Chawla (2025); Morales-Alvarado et al. (2024); Soelistyo and Lowe (2024); Tsoi et al. (2024). In this work, we use PySR Cranmer et al. (2020); Cranmer (2023), an open source SR library that effectively combines genetic algorithms with gradient-based optimization to learn symbolic models. This package has been widely applied in both astronomy and cosmology to find analytical expressions hidden in data, such as modeling the distribution of neutral hydrogen Wadekar et al. (2020), finding universal relations in subhalo properties Shao et al. (2022b), searching the galaxy-halo connection Delgado et al. (2022), reducing the Sunyaev-Zeldovich flux-mass scatter Wadekar et al. (2022), exploring feedback from active galactic nuclei Wadekar et al. (2023), discovering black hole (BH) mass scaling relations Jin and Davis (2023), rediscovering orbital mechanics Lemos et al. (2023), investigating the dynamical dark energy in DESI dataset Sousa-Neto et al. (2025) and so on.
@@
Here, we apply SR to the study of Gray-Body Factors (GBFs) in Primordial Black Holes (PBHs), which are compelling Dark Matter (DM) candidates Carr and Kuhnel (2020); Bird et al. (2023); Green (2024). Despite decades of experimental effort, the nature of DM remains unknown Bertone et al. (2005); Feng (2010); Cirelli et al. (2024). Several well-motivated candidates have been proposed, such as weakly interacting massive particles Roszkowski et al. (2018); Yuan et al. (2022a), axion/axion-like particles Adams et al. (2022); Yuan et al. (2021), fuzzy dark matter and ultralight dark matter Hu et al. (2000); Yuan et al. (2022b). Among these possibilities, PBHs are unique since their origin does not require beyond-Standard Model physics: they can form naturally in the early Universe from the collapse of large density fluctuations Hawking (1971); Carr et al. (2016a); Yuan et al. (2024). PBHs are unique in that they may span an extremely broad mass range, $10^{-5}\sim 10^{55}$arr and Hawking (1974); Carr (1975), distinguishing them from BHs formed through stellar evolution. A key theoretical ingredient in modeling PBHs evaporation are the GBFs, which quantify the frequency-dependent transmission probabilities of particles escaping the BH’s gravitational potential Dubinsky and Zinhailo (2024). These GBFs modulate the otherwise blackbody Hawking spectrum and are essential for accurate predictions of observational signatures. However, computing GBFs involves solving complex differential equations for each particle species and spin in curved spacetime, which is typically a numerically intensive task. To address this, we employ SR to derive analytic approximations for the numerically computed GBFs, using the recently developed GrayHawk package Calzá (2025). Our symbolic models not only reproduce the numerical results with high accuracy, but also offer compact, interpretable expressions suitable for fast evaluation. We demonstrate that these SR-based approximations can be used to place constraints on $f_{\rm PBH}$, the fraction of DM composed of PBHs, using measurements of the diffuse Extragalactic $\gamma$-Gay Background (EGRB). By comparing SR and numerical results, we validate the effectiveness of the former in reproducing constraints on $f_{\rm PBH}$.
@@
This approach is particularly relevant for assessing PBHs as DM candidates. A distinctive observational signature of low-mass PBHs is their Hawking radiation, a quantum mechanical effect by which BHs lose mass through the emission of a quasi-thermal spectrum of particles Hawking (1975a). This radiation encompasses photons Arbey et al. (2020); Laha (2019); Ballesteros et al. (2020), electrons/positrons MacGibbon and Carr (1991); Boudaud and Cirelli (2019); Su et al. (2024), neutrinos Dasgupta et al. (2020); Wang et al. (2021); Calabrese et al. (2022), and other Standard Model particles Carr et al. (2010); Clark et al. (2017); Acharya and Khatri (2020); Calzà and Rosa (2023); Calzà et al. (2024a, b) and beyond. This radiation emitted by PBH is expected to produce observable effects that could be identified in cosmic and astrophysical observations Belotsky et al. (2014); Carr et al. (2021); Carr and Kuhnel (2020); Auffinger (2023). More precisely, the emission from PBH accumulated across cosmic history should leave a detectable signature in $\gamma$-ray surveys Carr et al. (2016b); Chen et al. (2022). Analyzing the EGRB is therefore important for assessing how much of the DM may be comprised of PBHs. Our SR-based approximation allows for efficient and accurate calculation of the radiation predicted from PBH, thereby aiding in setting constraints on $f_{\rm PBH}$.
@@
This paper is organized as follows. Section II covers a review of Hawking radiation alongside the significance of GBFs and introduces the GrayHawk code employed for numerical computation. In Section III.1, we explain the principle and methodology of PySR, and describe its application to the fitting of GBFs, which is referred to as ReGrayssion herein, and we plan to make it public shortly. Section IV presents our approach to deriving constraints on $f_{\rm PBH}$ from $\gamma$-ray observations. Finally, we summarize the implications of our findings in Section V. Throughout this work, we adopt natural units $G=c=\hbar=k_{B}=1$.
@@ II.1 Hawking Spectrum
where $n_{i}$ is the number of degrees of freedom of the emitted particle, $\omega=E_{i}$ is the mode frequency, and the GBFs $\Gamma^{s}_{l}(\omega)$ determine the transmission probability for each mode. The sign at the denominator accounts for the Bose-Einstein or Fermi-Dirac statistics, -- and $+$ respectively, while the factor $(2l+1)$ in Eq. (5) arises from the degeneracy in the azimuthal number $m$ due to spherical symmetry. Eq. (5) encapsulates the key role of GBFs in shaping the energy spectrum of Hawking radiation, making their accurate computation crucial for studying BH evaporation.
@@ III.1 PySR
PySR employs a multi-population evolutionary algorithm that systematically searches for candidate equations from a predefined set of mathematical operators and variables. Each symbolic expression is assessed based on a scoring mechanism that incorporates both accuracy and complexity; the latter can be regulated by a user-defined parsimony coefficient. Accuracy is typically evaluated via a loss function, while complexity is quantified as a weighted sum of the number of variables, operators, and constants in the expression. Table 1 lists the binary operators (+,,×,÷$+,-,\times,\div,{\rm pow}$) and unary operators (sin, cos, exp, log, sqrt, erf, erfc) adopted in this work.
@@ III.3 Methodology: GBF analytic approximations
By default, variable and parameter complexities are set to 1, making the complexity of the polynomial ($a+b*l+c*\omega$) 10. Thus, the total complexity of Eq. (10) becomes 18 with the addition of ‘erf’ and ‘erfc’(see Fig. 3). Although we remove the ‘erfc’ and add two operators (×,$\times$,+× , +) and parameters, increasing the complexity to 19, Eq. (11) remains more interpretable than Eq. (10).
@@ IV Constraining PBH as DM Using Analytical GBFs
The possibility of PBH accounting for the entire DM budget is highly limited by numerous observations Katz et al. (2018); Bai and Orlofsky (2019); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2024); Gorton and Green (2024); Dent et al. (2025); Tamta et al. (2025); Tinyakov (2024); Loeb (2024). The only remaining parameter space where PBHs might constitute all DM is the asteroid mass window, approximately spanning PBH masses from $10^{17}{\rm g}$ to $10^{23}{\rm g}$ Montero-Camacho et al. (2019); Coogan et al. (2021); Auffinger (2023); Dent et al. (2025). PBHs with lower masses would either have vanished due to Hawking evaporation or emitted too many MeV $\gamma$-rays to be consistent with the observations.
@@ IV.1 Photon Spectrum from PBH Evaporation
where $\rho_{\rm crit,0}=3H_{0}^{2}/8\pi G$ is the critical density. We assume a $\Lambda$CDM cosmology with cosmological parameters from the Planck experiment Aghanim et al. (2020). We notice that the findings remain consistent with reasonable variations in these parameters. This framework provides a precise evaluation of PBH abundance constraints from Hawking evaporation and EGRB observations, further discussed in Sec. IV.2.
@@ IV.2 Observational Constraints on $f_{\rm PBH}$
$\Gamma^{s=1}_{l}(\omega)=\begin{cases}{\rm erf}[{\rm erfc}(a+bl+c\omega)]\quad{\rm SR1},\\ {\rm erf}[{\rm erfc}(a+bl+c\omega+d\omega^{2})]\quad{\rm SR2},\\ {\rm erf}[{\rm erfc}(a+bl+c\omega+d\omega^{2}+e\omega^{3})]\quad{\rm SR3}.\\ \end{cases}$ (15)
@@ Appendix A Fitting GBFs with Symbolic Regression
Figure 6: The four subplots depicting various fits of $\Gamma(\omega,l)$ using symbolic regression (SR), each subjected to different input error rates: $\epsilon=10\%,5\%,1\%$, and $0.1\%$ as described in Eq. (8). Data points embedded with errors are plotted, alongside lines generated by functions indicated in the subplots’ titles. The colors blue, orange, green, and red correspond to cases where $l=1,2,3,4$, respectively. The figure illustrates that as the magnitude of errors increases, it becomes increasingly challenging for the SR technique to derive the analytic expressions.

### v3

@@ I Introduction
By defining a target variable, a collection of potential explanatory variables, and a set of available mathematical operators, SR aims to find equations that accurately predict the target variable as a function of the explanatory ones, while remaining concise and interpretable Makke and Chawla (2024); Landajuela et al. (2022); Wang et al. (2023). Various SR frameworks are currently accessible, such as EUREQA Schmidt and Lipson (2009), Operon Burlacu et al. (2020), DSR Petersen et al. (2021), EQL Sahoo et al. (2018), ESR Bartlett et al. (2024a), and AI Feynman Udrescu and Tegmark (2020). These interpretable machine learning tools have recently sparked renewed interest, catalyzing notable progress across both theoretical and applied sciences La Cava et al. (2021); Bernal et al. (2021); Butter et al. (2024); Shao et al. (2022a); Li et al. (2023); Bartlett et al. (2023); Angelis et al. (2023); Koksbang (2023a, b); Desmond et al. (2023); Sousa et al. (2024); Bartlett et al. (2024b, 2023); Koksbang (2023c); Thing and Koksbang (2025); Bartlett et al. (2024c); Constantin et al. (2024); Sui et al. (2024); Makke and Chawla (2025); Morales-Alvarado et al. (2024); Soelistyo and Lowe (2024); Tsoi et al. (2024). In this work, we use PySR Cranmer et al. (2020); Cranmer (2023), an open source SR library that effectively combines genetic algorithms with gradient-based optimization to learn symbolic models. This package has been widely applied in both astronomy and cosmology to find analytical expressions hidden in data, such as modeling the distribution of neutral hydrogen Wadekar et al. (2020), finding universal relations in subhalo properties Shao et al. (2022b), searching the galaxy-halo connection Delgado et al. (2022), reducing the Sunyaev-Zeldovich flux-mass scatter Wadekar et al. (2022), exploring feedback from active galactic nuclei Wadekar et al. (2023), discovering black hole (BH) mass scaling relations Jin and Davis (2023), rediscovering orbital mechanics Lemos et al. (2023), investigating the dynamical dark energy Sousa-Neto et al. (2025) and so on.
@@
Here, we apply SR to the study of Gray-Body Factors (GBFs) in Primordial Black Holes (PBHs), which are compelling Dark Matter (DM) candidates Carr and Kuhnel (2020); Bird et al. (2023); Green (2024). Despite decades of experimental effort, the nature of DM remains unknown Bertone et al. (2005); Feng (2010); Cirelli et al. (2024). Several well-motivated candidates have been proposed, such as weakly interacting massive particles Roszkowski et al. (2018); Yuan et al. (2022a), axion/axion-like particles Adams et al. (2022); Yuan et al. (2021), fuzzy dark matter and ultralight dark matter Hu et al. (2000); Yuan et al. (2022b). Among these possibilities, PBHs are unique since their origin does not require beyond-Standard Model physics: they can form naturally in the early Universe from the collapse of large density fluctuations Hawking (1971); Carr et al. (2016a); Yuan et al. (2024). PBHs are unique in that they may span an extremely broad mass range, $10^{-5}\sim 10^{55}$Carr and Hawking (1974); Carr (1975), distinguishing them from BHs formed through stellar evolution. A key theoretical ingredient in modeling PBHs evaporation is the GBFs, which quantify the frequency-dependent transmission probabilities of particles escaping the BH’s gravitational potential Dubinsky and Zinhailo (2024). These GBFs modulate the otherwise blackbody Hawking spectrum and are essential for accurate predictions of observational signatures. However, computing GBFs involves solving complex differential equations for each particle species and spin in curved spacetime, which is typically a numerically intensive task. To address this, we employ SR to derive analytic approximations for the numerically computed GBFs, using the recently developed GrayHawk package Calzá (2025). Our symbolic models not only reproduce the numerical results with high accuracy, but also offer compact, interpretable expressions suitable for fast evaluation. We demonstrate that these SR-based approximations can be used to place constraints on $f_{\rm PBH}$, the fraction of DM composed of PBHs, using measurements of the diffuse Extragalactic $\gamma$-Gay Background (EGRB). By comparing SR and numerical results, we validate the effectiveness of the former in reproducing constraints on $f_{\rm PBH}$.
@@
This approach is particularly relevant for assessing PBHs as DM candidates. A distinctive observational signature of low-mass PBHs is their Hawking radiation, a quantum mechanical effect by which BHs lose mass through the emission of a quasi-thermal spectrum of particles Hawking (1975a); Saini and Stojkovic (2018). This radiation encompasses photons Arbey et al. (2020); Laha (2019); Ballesteros et al. (2020); Cang et al. (2022); Perez-Gonzalez (2023), electrons/positrons MacGibbon and Carr (1991); Boudaud and Cirelli (2019); Su et al. (2024), neutrinos Dasgupta et al. (2020); Wang et al. (2021); Lunardini and Perez-Gonzalez (2020); Calabrese et al. (2022); De Romeri et al. (2025), and other Standard Model particles Stojkovic and Freese (2005); Carr et al. (2010); Clark et al. (2017); Acharya and Khatri (2020); Calzà and Rosa (2023); Calzà et al. (2024a, b) and beyond Stojkovic et al. (2005); Dong et al. (2016); Cang et al. (2023). This radiation emitted by PBH is expected to produce observable effects that could be identified in cosmic and astrophysical observations Belotsky et al. (2014); Carr et al. (2021); Carr and Kuhnel (2020); Auffinger (2023). More precisely, the emission from PBH accumulated across cosmic history should leave a detectable signature in $\gamma$-ray surveys Carr et al. (2016b); Chen et al. (2022). Analyzing the EGRB is therefore important for assessing how much of the DM may be comprised of PBHs. Our SR-based approximation allows for efficient and accurate calculation of the radiation predicted from PBH, thereby aiding in setting constraints on $f_{\rm PBH}$.
@@
This paper is organized as follows. Section II covers a review of Hawking radiation alongside the significance of GBFs and introduces the GrayHawk code employed for numerical computation. In Section III.1, we explain the principle and methodology of PySR, and describe its application to the fitting of GBFs, which is referred to as ReGrayssion herein. Additionally, it can be accessed publicly on GitHub through the link: https://github.com/yuanguanwen/Regrayssion. Section IV presents our approach to deriving constraints on $f_{\rm PBH}$ from $\gamma$-ray observations. Finally, we summarize the implications of our findings in Section V. Throughout this work, we adopt natural units $G=c=\hbar=k_{B}=1$.
@@ II.1 Hawking Spectrum
where $n_{i}$ is the number of degrees of freedom of the emitted particle, $\omega=E_{i}$ is the mode frequency, and the GBFs $\Gamma^{s}_{l}(\omega)$ determine the transmission probability for each mode. The sign at the denominator accounts for the Bose-Einstein or Fermi-Dirac statistics, - and ++ respectively, while the factor $(2l+1)$ in Eq. (5) arises from the degeneracy in the azimuthal number $m$ due to spherical symmetry. Eq. (5) encapsulates the key role of GBFs in shaping the energy spectrum of Hawking radiation, making their accurate computation crucial for studying BH evaporation.
@@ III.1 PySR
PySR employs a multi-population evolutionary algorithm that systematically searches for candidate equations from a predefined set of mathematical operators and variables. Each symbolic expression is assessed based on a scoring mechanism that incorporates both accuracy and complexity; the latter can be regulated by a user-defined parsimony coefficient. Accuracy is typically evaluated via a loss function, while complexity is quantified as a weighted sum of the number of variables, operators, and constants in the expression. Table 1 lists the binary operators ($\times,\div,{\rm pow}$) and unary operators (sin, cos, exp, log, sqrt, erf, erfc) adopted in this work.
@@ III.3 Methodology: GBF analytic approximations
By default, variable and parameter complexities are set to 1, making the complexity of the polynomial ($a+b*l+c*\omega$) 10. Thus, the total complexity of Eq. (10) becomes 18 with the addition of ‘erf’ and ‘erfc’(see Fig. 3). Although we remove the ‘erfc’ and add two operators ($\times,+$) and parameters, increasing the complexity to 19, Eq. (11) remains more interpretable than Eq. (10).
@@ IV Constraining PBH as DM Using Analytical GBFs
The possibility of PBH accounting for the entire DM budget is highly limited by numerous observations Katz et al. (2018); Bai and Orlofsky (2019); Smyth et al. (2020); Coogan et al. (2021); Ray et al. (2021); Auffinger (2022); Ghosh and Mishra (2024); Miller et al. (2022); Branco et al. (2023); Bertrand et al. (2023); Tran et al. (2024); Gorton and Green (2024); Dent et al. (2025); Tamta et al. (2025); Tinyakov (2024); Loeb (2024). The only remaining parameter space where PBHs might constitute all DM is the asteroid mass window, approximately spanning PBH masses from $10^{17}{\rm g}$ to $10^{23}{\rm g}$ Montero-Camacho et al. (2019); Coogan et al. (2021); Auffinger (2023); Dent et al. (2025); Dai and Stojkovic (2024). PBHs with lower masses would either have vanished due to Hawking evaporation or emitted too many MeV $\gamma$-rays to be consistent with the observations.
@@ IV.1 Photon Spectrum from PBH Evaporation
[…]
@@ IV.2 Observational Constraints on $f_{\rm PBH}$
[…]
@@ Appendix A Fitting GBFs with Symbolic Regression
Figure 6: The four subplots depicting various fits of $\Gamma(\omega,l)$ using symbolic regression (SR), each subjected to different input error rates: $\epsilon=10\%,5\%,1\%$, and $0.1\%$ as described in Eq. (8). Data points embedded with errors are plotted, alongside lines generated by functions indicated in the subplots’ titles. The colors purple, blue, orange and green correspond to cases where $l=1,2,3,4$, respectively. The figure illustrates that as the magnitude of errors increases, it becomes increasingly challenging for the SR technique to derive the analytic expressions.